OpenStudy (howard-wolowitz):

Triangle question:

OpenStudy (anonymous):

\[cosec(B) = \frac{Hypotenuse}{Perpendicular} = \frac{c}b{}\]

OpenStudy (howard-wolowitz):

where do u get c

OpenStudy (anonymous):

cosec(B) = 10 b = 0.5

OpenStudy (welshfella):

and you know that csc B = 10 = c /b and b = 0.5

OpenStudy (anonymous):

c is Hypotenuse there. So I replaced Hypotenuse with c there.

OpenStudy (howard-wolowitz):

alright gotcha

OpenStudy (welshfella):

plug in b = 0.5 10 = c / 0.5 so now you need to solve for c

OpenStudy (anonymous):

\[10 = \frac{c}{0.5}\]

OpenStudy (anonymous):

Find \(c\) from here now..

OpenStudy (howard-wolowitz):

its 5

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

you are getting more confident Howard

OpenStudy (howard-wolowitz):

I"m trying to understand more then before